OpenStudy (anonymous):

hey guys, sorry to bother u but can someone show me how to solve this equation with the steps included ? thanks :) 2 (k+5) -3 (k-1)= k - 7

OpenStudy (anonymous):

It's just a matter of "collecting" all the k terms...

OpenStudy (anonymous):

oops, algebra error in my last post... you just want to multiply everything out, then simplify: So you just want to multiply everything out and then simplify: 2 (k+5) -3 (k-1) = k - 7 2k + 10 -3k +3 = k - 7 Then collect the k terms on the left side and the number terms on the right: 2k + 10 -3k +3 = k - 7 2k - 3k - k = -7 - 10 - 3 then simplify 2k - 3k - k = -7 - 10 - 3 -2k = -20 Then divide both sides by -2: -2k/(-2) = -20/(-2) Finally, simplify to solve for k: -2k/(-2) = -20/(-2) -2k/(-2) = 10 k = 10

OpenStudy (anonymous):

So, @helloxox, that's a pretty detailed solution... please read through the process and let me know if you have questions, 'cause I will be happy to explain further. Hopefully after reading it, you get the process of step by step simplifying an equation to isolate or "solve for" the variable.

OpenStudy (anonymous):

thank u so much :))

OpenStudy (anonymous):

Does it make sense now? Hope so... :)

OpenStudy (hba):

@JakeV8 Please Don't Answer Questions Directly :)

OpenStudy (anonymous):

ok ?

OpenStudy (hba):

I Think you should check the CoC- http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Sometimes seeing an example worked is helpful.... if the asker, like @helloxox, is one of those who just posts "SOLVE!", then I will be sure not to help further... In an asynchronous situation like this, sometimes an example helps get the conversation started.

OpenStudy (hba):

It would be better if you would involve the user to participate in the question.

OpenStudy (anonymous):

Trust me, I'm not handing out answers repeatedly, but seeing a problem worked, with explanation included, may well help a student who wants to learn. Someone who is answer-trolling will take it and run... I get that... but I will remember the lack of effort and not reward it twice, and I will be happy to DM other helpers to give them a heads up on the issue.

OpenStudy (anonymous):

@sbanuelos, so @hba is talking about you here... make sure you're learning, not just capturing answers from people like me who are trying to help teach.

OpenStudy (hba):

I hope @sbanuelos is learning and not in search of answers like other users :) Anyways,@JakeV8 It Would be better if you would solve a sample question for them,which looks quiet similar with explanation.